OpenStudy (anonymous):

A and B are two matrices ,such that AB = A and BA = B. then what is the value of B square ?

OpenStudy (unklerhaukus):

do they both have to be the identity matrix ?

OpenStudy (anonymous):

i can't say so. i think maybe it is.

OpenStudy (unklerhaukus):

\[\textbf{AB} = \textbf A\]dividing on the left by \(\textbf{A}\)\[\textbf{B}=\textbf{I}\] and \[\textbf{BA} = \textbf{B}\]dividing on the left by \(\textbf B\)\[\textbf A=\textbf I\]

OpenStudy (unklerhaukus):

\[\textbf B^2=\]

OpenStudy (anonymous):

we need to find the value of \[{B}2\]

OpenStudy (unklerhaukus):

Well i just proved \[\textbf B=\textbf I\] \[\textbf B^2=\textbf I^2=\]

OpenStudy (unklerhaukus):

am i making sense ?

OpenStudy (anonymous):

how is that possible.can you please elaborate ? @UnkleRhaukus

OpenStudy (unklerhaukus):

\[\textbf {II}=\textbf I\]

OpenStudy (unklerhaukus):

where is the confusion

OpenStudy (anonymous):

thanks @UnkleRhaukus